# Evolutionary Prediction Games

### v1

@@ B.5.1 Data distribution
B.5.2 Stable beneficial coexistence
We show that the Soft-SVM algorithm can induce a stable, beneficial coexistence. We will consider learning in the population limit $n\to\infty$, where the loss minimization objective (Definition B.15) converges towards its expected value:
Definition B.16 (Soft-SVM in the population limit).
@@ B.5.2 Stable beneficial coexistence
Proof of Theorem 5.1.
Assume the Soft-SVM algorithm trains on a large dataset sampled from $D_{\bm{p}}$. Denote ${\bm{p}}=(1-p,p)$, and denote the probability density functions of groups ${\left\{{A,B}\right\}}$ by $f_{A}(x,y),f_{B}(x,y)$, respectively. Since the data is separable and labels are positively correlated with the feature $x$, it holds that $w^{*}>0$. Assume that regularization parameter $\gamma$ is large enough such that $w^{*}\leq 1$. From Proposition B.17, it holds that:
@@
When $w\in(0,1]$, observe that the distribution defined in Section B.5.1 is suppported on $[-1,1]$, and therefore both integrals cover the whole distribution. We can therefore write:
@@ B.5.3 The effect of regularization
In Figure 7, we vary the regularization strength for the distributions specified in Section B.5.1 with imbalance parameter $\alpha=0.75$. We observe that the system transition between three topologically-distinct phases: For weak regularization ($\lambda=10$), coexistence is beneficial but not stable. For intermediate regularization ($\lambda=100$), coexistence is beneficial and stable. Interestingly, for strong regularization ($\lambda=1000$), the game has five equilibria – Two stable single-group equilibria, one stable coexistence equilibia, and two unstable coexistence equilibria.
B.6 k-NN Classifiers With Label Noise
Definition B.18 (k-Nearest-Neighbors (k-NN); e.g. Shalev-Shwartz & Ben-David (2014), Section 19.1).
@@ B.6.1 Data distribution
B.6.2 Coexistence
Proposition B.19 (Convex combinations).
@@ B.6.2 Coexistence
Plugging Equation 11 into Equation 10, we obtain that the corresponding $q$ for the positive mixture is:
@@
Proposition B.20 (Reflection-exchange symmetry).
@@
Proposition B.21 (Expected prediction of $k$-NN).
@@
Lemma B.22 (Expected accuracy).
@@
Where the identity $\mathbb{E}{\left[{y}\right]}=(1-2\alpha)$ is given by the definition of $\tilde{D}_{\mathrm{pos}}^{\alpha}$, the definition of $\gamma$ is given by Proposition B.19, and the identity $\mathbb{E}{\left[{y_{i^{*}(x)}}\right]}=\left(-1+2\phi_{k}\left(\frac{\beta(1-p)}{\gamma}\alpha\right)\right)$ is given by Proposition B.21.
@@
Figure 8: k-NN accuracy for the construction specified in Section B.6.1. The plots show alignment between the theoretical curves (Lemma B.22) and empirical simulations (mean accuracies represented by dots). For $k=3,5$ the game has a stable beneficial coexistence.
Proposition B.23.
@@
By applying Lemma B.22. When $k=1$, the $\phi_{k}$ function satisfies:
@@
Plugging into Equation 12 and using the symmetry argument in Proposition B.20, we obtain:
@@
Proof of Theorem 5.2.
By Proposition B.23, using e.g. $(\alpha,\beta)=(0.2,0.8)$. We observe that the conditions of the proposition are satisfied for this choice of constants, as $\alpha=0.2<0.375=1-\frac{1}{2\cdot 0.8}=1-\frac{1}{2\beta}$. ∎
B.7 Hard-SVM With Finite Data
Definition B.24 (Hard-SVM; e.g. Shalev-Shwartz & Ben-David (2014), Section 15.1).
@@ B.7 Hard-SVM With Finite Data
Remark B.25 (One-dimensional Hard-SVM).
@@ B.7.2 Coexistence
By Equation 15 we obtain the expectation of $x_{\mathrm{min},n}$:
@@
Equation 14 gives a lower bound on the expectation:
@@
For case (1), the Hard-SVM algorithm trains on data sampled from $D_{A}$. Denote the size of the dataset by $n$ and the corresponding Hard-SVM classfier by $h_{n}$. We note that $h_{n}$ is a random variable depending on the dataset. For a dataset ${\left\{{(x_{i},y_{i})}\right\}}_{i=1}^{n}\sim D_{A}^{n}$, Since the problem is one-dimensional, the Hard-SVM decision margin is given by Equation 13:
@@
By Proposition B.26 and for $n\geq 15$, it holds that $\mathbb{E}{\left[{x_{\mathrm{margin}}}\right]}\geq\frac{1}{n}+\varepsilon$, and therefore:
@@
Proof of Theorem 5.3.
B.8 Stabilizing Coexistence
Proof of Proposition 5.4.
Let $F({\bm{p}})$ be an evolutionary prediction game induced by an optimal learning algorithm $\mathcal{A}({\bm{p}})$, and let ${\bm{p}}^{*}$ be an unstable coexistence equilibrium with full support (${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}=K$). The game induced by $\mathcal{A}^{\prime}({\bm{p}})$ in the neighborhood of ${\bm{p}}^{*}$ is $F^{\prime}({\bm{p}})=F(2{\bm{p}}^{*}-{\bm{p}})$. By Lemma B.14, the game $F({\bm{p}})$ is a potential game, and therefore admits a potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$. Consider the potential function $f^{\prime}({\bm{p}})=-f(2{\bm{p}}^{*}-{\bm{p}})$. By the chain rule, it holds that:
@@ B.8 Stabilizing Coexistence
And therefore $f^{\prime}({\bm{p}})$ is a potential function for the game $F^{\prime}({\bm{p}})$ in the neighborhood of ${\bm{p}}^{*}$. The equilibrium ${\bm{p}}^{*}$ is unstable, and therefore by (Sandholm, 2010, Theorem 8.2.1) it is not a maximizer of $f({\bm{p}})$. Moreover, by Proposition B.9 the potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex, and as ${\bm{p}}^{*}$ is assumed to have full support, it is also a global minimizer of $f({\bm{p}})$. From this we conclude that ${\bm{p}}^{*}$ is a global maximizer of $f^{\prime}({\bm{p}})$, and therefore the equilibrium ${\bm{p}}^{*}$ is stable under the evolutionary game induced by $\mathcal{A}^{\prime}$. ∎
Appendix C Experimental Details
@@
We implement our analysis in Python. Our synthetic-data experiments rely on scikit-learn for learning algorithm implementations (Pedregosa et al., 2011), and our real-data experiments rely on PyTorch (Paszke et al., 2019) and ffcv (Leclerc et al., 2023). Code is available at: https://github.com/edensaig/evolutionary-prediction-games.
@@
Synthetic data simulations were run on a single Macbook Pro laptop, with 16GB of RAM, M2 processor, and no GPU support. Experiments involving neural networks (Section 6) were run on a dedicated server with an AMD EPYC 7502 CPU, 503GB of RAM, and an Nvidia RTX A4000 GPU.
Evolutionary dynamics.
For simulation of evolutionary dynamics (Figure 3, Figure 5), we use the discrete replicator equation given by (Taylor & Jonker, 1978, Equation (3)).
Architectures.
For CIFAR-10, we use the Resnet-9 architecture and training code provided in the CIFAR-10 example code in the ffcv Github repository (libffcv/ffcv, commit 7885f40), with modifications to control the probability of horizontal flips in training and testing. For MNIST, we use the convolutional neural network provided in the MNIST example code in the PyTorch examples Github repository (pytorch/examples, commit 37a1866). The MNIST network has two convolutional layers, and two fully-connected layers, with dropout and max pooling. Training is performed for 200 epochs using SGD with learning rate $0.01$ and momentum $0.5$. We use ffcv for data loading in both experiments.

### v2

@@ D.6.1 Construction
D.6.2 Stable Mutualistic Coexistence
We show that the Soft-SVM algorithm can induce a stable, beneficial coexistence. We will consider learning in the population limit $n\to\infty$, where the loss minimization objective (Definition 10) converges towards its expected value:
Definition 11 (Soft-SVM in the population limit).
@@ D.6.2 Stable Mutualistic Coexistence
Proof of Theorem 1.
Assume the Soft-SVM algorithm trains on a large dataset sampled from $D_{\bm{p}}$. Denote ${\bm{p}}=(1-p,p)$, and denote the probability density functions of groups ${\left\{{A,B}\right\}}$ by $f_{A}(x,y),f_{B}(x,y)$, respectively. Since the data is separable and labels are positively correlated with the feature $x$, it holds that $w^{*}>0$. Assume that regularization parameter $\gamma$ is large enough such that $w^{*}\leq 1$. From Proposition 14, it holds that:
@@
When $w\in(0,1]$, observe that the distribution defined in Section D.6.1 is suppported on $[-1,1]$, and therefore both integrals cover the whole distribution. We can therefore write:
@@ D.6.3 Bifurcations Induced by Regularization
In Figure 9, we vary the regularization strength for the distributions specified in Section D.6.1 with imbalance parameter $\alpha=0.75$. We observe that the system transition between three topologically-distinct phases: For weak regularization ($\lambda=10$), coexistence is beneficial but not stable. For intermediate regularization ($\lambda=100$), coexistence is beneficial and stable. Interestingly, for strong regularization ($\lambda=1000$), the game has five equilibria – Two stable single-group equilibria, one stable coexistence equilibia, and two unstable coexistence equilibria.
D.7 k-NN Classifiers With Label Noise
Another practical approach to loss minimization is to use a class of models that fit the training data perfectly and then interpolate. Notable examples include the nearest-neighbor algorithm, and over-parameterized neural networks [e.g., 54]. Interpolation leads to perfect accuracy on the training set by definition, but often at the price of sensitivity to noise; as a concrete example, consider how 1-NN predictions change with the addition of a single point. The problem is that even in the limit, 1-NN cannot express the true $p(y|x)$, and remains sensitive to label noise. Our next result shows this can lead to coexistence:
Theorem 4.
Formal proof in Section D.7.2. The construction is illustrated in Figure 3 (Center), and described formally in Section D.7.1. Informally, there are two groups defined symmetrically. Each group is composed of a majority and minority classes, and the majority class has label noise. For each group alone, the 1-NN algorithm makes perfect predictions on its minority class, and imperfect predictions on the majority class. Since groups are defined symmetrically, training on one group alone leads to better accuracy for the other—leading to stable coexistence. Moreover, Lemma 2 shows that for $k$-NN with $k\geq 3$, stable coexistence can also be beneficial for both groups.
Definition 12 (k-Nearest-Neighbors; e.g. [45, Section 19.1]).
@@ D.7.1 Construction
D.7.2 Stable Coexistence
Proposition 15 (Convex combinations).
@@ D.7.2 Stable Coexistence
Plugging eq. 12 into eq. 11, we obtain that the corresponding $q$ for the positive mixture is:
@@
Proposition 16 (Reflection-exchange symmetry).
@@
Proposition 17 (Expected prediction of $k$-NN).
@@
Lemma 2 (Expected accuracy).
@@
Where the identity $\mathbb{E}{\left[{y}\right]}=(1-2\alpha)$ is given by the definition of $\tilde{D}_{\mathrm{pos}}^{\alpha}$, the definition of $\gamma$ is given by Proposition 15, and the identity $\mathbb{E}{\left[{y_{i^{*}(x)}}\right]}=\left(-1+2\phi_{k}\left(\frac{\beta(1-p)}{\gamma}\alpha\right)\right)$ is given by Proposition 17.
@@
Figure 10: k-NN accuracy for the construction specified in Section D.7.1. The plots show alignment between the theoretical curves (Lemma 2) and empirical simulations (mean accuracies represented by dots). For $k=3,5$ the game has a stable beneficial coexistence.
Proposition 18.
@@
By applying Lemma 2. When $k=1$, the $\phi_{k}$ function satisfies:
@@
Plugging into eq. 13 and using the symmetry argument in Proposition 16, we obtain:
@@
Proof of Theorem 4.
By Proposition 18, using e.g. $(\alpha,\beta)=(0.2,0.8)$. We observe that the conditions of the proposition are satisfied for this choice of constants, as $\alpha=0.2<0.375=1-\frac{1}{2\cdot 0.8}=1-\frac{1}{2\beta}$. ∎
D.8 Hard-SVM With Finite Data
All practical learning algorithms must rely on a finite sample of data points for training a classifier. Learnability (in the PAC sense) implies that as data size grows, the sample becomes increasingly representative of the distribution. But the rate at which variance diminishes need not be equally across all regions of a distribution. In a sense, this means that one region can be learned ‘faster’ (i.e., with less data) than another. Our final example exploits this to enable coexistence.
Theorem 5.
Proof in Section D.8.2. The construction is illustrated in Figure 3 (Right), and described formally in Section D.8.1. Informally, the key property we leverage here is asymmetric class variance. In each group, one class has high variance, and the other class has low variance. When the dataset is of size $n$, this creates bias against the high-variance class, and thus convergence towards the optimal classifier at rate $O(n^{-1})$. When the two groups coexist, this bias is balanced, and the rate of convergence is exponential—leading to coexistence that is beneficial, but unstable.
Definition 13 (Hard-SVM; e.g. [45, Section 15.1]).
@@ D.8 Hard-SVM With Finite Data
Remark 1 (One-dimensional Hard-SVM).
@@ D.8.2 Mutualistic Coexistence
By eq. 16 we obtain the expectation of $x_{\mathrm{min},n}$:
@@
eq. 15 gives a lower bound on the expectation:
@@
For case (1), the Hard-SVM algorithm trains on data sampled from $D_{A}$. Denote the size of the dataset by $n$ and the corresponding Hard-SVM classfier by $h_{n}$. We note that $h_{n}$ is a random variable depending on the dataset. For a dataset ${\left\{{(x_{i},y_{i})}\right\}}_{i=1}^{n}\sim D_{A}^{n}$, Since the problem is one-dimensional, the Hard-SVM decision margin is given by eq. 14:
@@
By Proposition 19 and for $n\geq 15$, it holds that $\mathbb{E}{\left[{x_{\mathrm{margin}}}\right]}\geq\frac{1}{n}+\varepsilon$, and therefore:
@@
Proof of Theorem 5.
By Proposition 20. ∎
D.9 Stabilizing Coexistence
Proof of Proposition 1.
Let $F({\bm{p}})$ be an evolutionary prediction game induced by an optimal learning algorithm $\mathcal{A}({\bm{p}})$, and let ${\bm{p}}^{*}$ be an unstable coexistence equilibrium with full support (${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}=K$). The game induced by $\mathcal{A}^{\prime}({\bm{p}})$ in the neighborhood of ${\bm{p}}^{*}$ is $F^{\prime}({\bm{p}})=F(2{\bm{p}}^{*}-{\bm{p}})$. By Lemma 1, the game $F({\bm{p}})$ is a potential game, and therefore admits a potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$. Consider the potential function $f^{\prime}({\bm{p}})=-f(2{\bm{p}}^{*}-{\bm{p}})$. By the chain rule, it holds that:
@@ D.9 Stabilizing Coexistence
And therefore $f^{\prime}({\bm{p}})$ is a potential function for the game $F^{\prime}({\bm{p}})$ in the neighborhood of ${\bm{p}}^{*}$. The equilibrium ${\bm{p}}^{*}$ is unstable, and therefore by [43, Theorem 8.2.1] it is not a maximizer of $f({\bm{p}})$. Moreover, by Proposition 3 the potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex, and as ${\bm{p}}^{*}$ is assumed to have full support, it is also a global minimizer of $f({\bm{p}})$. From this we conclude that ${\bm{p}}^{*}$ is a global maximizer of $f^{\prime}({\bm{p}})$, and therefore the equilibrium ${\bm{p}}^{*}$ is stable under the evolutionary game induced by $\mathcal{A}^{\prime}$. ∎
Appendix E Experiment Details
@@
We implement our analysis in Python. Our synthetic-data experiments rely on scikit-learn for learning algorithm implementations [38], and our real-data experiments rely on PyTorch [37] and ffcv [29]. We use the matplotlib package for plotting, and the mpltern package for ternary plots. Code is available at: https://github.com/edensaig/evolutionary-prediction-games.
@@
Synthetic data simulations were run on a single Macbook Pro laptop, with 16GB of RAM, M2 processor, and no GPU. Experiments involving neural networks (Section 6) were run on a dedicated server with an AMD EPYC 7502 CPU, 503GB of RAM, and an Nvidia RTX A4000 GPU.
@@
Architectures.
For CIFAR-10, we use the Resnet-9 architecture and training code provided in the CIFAR-10 example code in the ffcv Github repository (libffcv/ffcv, commit 7885f40), with modifications to control the probability of horizontal flips in training and testing. For MNIST, we use the convolutional neural network provided in the MNIST example code in the PyTorch examples Github repository (pytorch/examples, commit 37a1866). The MNIST network has two convolutional layers, and two fully-connected layers, with dropout and max pooling. Training is performed for 200 epochs using SGD with learning rate $0.01$ and momentum $0.5$.
Replicator dynamics simulation.
In the experiments, we use the discrete replicator equation for simulating evolutionary dynamics. The continuous-time replicator equation is:
$\dot{p}_{k}=p_{k}\left(F_{k}({\bm{p}})-\bar{F}({\bm{p}})\right)$
where $\bar{F}({\bm{p}})=\sum_{k}p_{k}F_{k}({\bm{p}})$ is the average fitness across the population. In our simulations (e.g., Figures 2, 4), we use the discrete-time replicator equation given by Taylor and Jonker [48, eq. (3)]:
$p^{t+1}_{k}=p^{t}_{k}\frac{F_{k}({\bm{p}})+1}{\bar{F}({\bm{p}})+1}$ (17)
For the CIFAR-10 replicator simulation (Section 6.2), we discretize the two-group simplex into the grid ${\left\{{(1,0),(0.9,0.1),\dots,(0,1)}\right\}}$, precompute fitness values $F_{k}({\bm{p}})$ at each grid point, and use linear interpolation to determine fitness at intermediate states.
Confidence intervals.
In both experiments, variability is due to the random split within the train and test sets, and due to random noise when it is added. For numerical results, we report normal mean confidence intervals at $99\%$ confidence level.